OpenStudy (anonymous):

verify csc^2x = cos^2x + cot^2x + sin^2x

hero (hero):

\(\csc^2x = \dfrac{1}{\sin^2x}\) Right @mgs122995 ?

OpenStudy (anonymous):

right!!

hero (hero):

\( \dfrac{1}{\sin^2x} = \dfrac{\sin^2x + \cos^2x}{\sin^2x}\) Right?

OpenStudy (anonymous):

yes

hero (hero):

\(\dfrac{\sin^2x + \cos^2x}{\sin^2x} = \dfrac{\sin^2x}{\sin^2x} + \dfrac{\cos^2x}{\sin^2x}\) Right?

OpenStudy (anonymous):

yes

hero (hero):

Do you have it from here or do you want me to keep going?

OpenStudy (anonymous):

keep going please

hero (hero):

\(\dfrac{\sin^2x}{\sin^2x} + \dfrac{\cos^2x}{\sin^2x} = 1 + \cot^2x\) right?

OpenStudy (anonymous):

okay yes

hero (hero):

\(1 + \cot^2x = (\sin^2x + \cos^2x) + \cot^2x\) right?

OpenStudy (anonymous):

yes

hero (hero):

Okay, so do you have any questions about how I got from the LHS to the RHS?

OpenStudy (anonymous):

no I understand that part

hero (hero):

Okay, great. Good luck with the rest of your work.

OpenStudy (anonymous):

thanks

hero (hero):

yw